OpenStudy (anonymous):

Please help me!!! ASAP!! 1. what is the solution of the matrix equation ? [9 4] x= [-9 -6] [2 1] [-1 -8] A. [5 -26] [9 -60] B. [-5 26] [-1 -8] C. [-5 26] [9 -60] D. [9 26] [9 1]

OpenStudy (kainui):

If you left multiply both sides of the equation by the inverse of the left matrix you will have the identity matrix on the left multiplied by X leaving just X behind so that will mean the right side will be your answer!

OpenStudy (anonymous):

oh okay so wait what do you multiply by?

OpenStudy (johnweldon1993):

\[\large \left[\begin{matrix}9 & 4 \\ 2 & 1\end{matrix}\right]x = \left[\begin{matrix}-9 & -6 \\ -1 & -8\end{matrix}\right]\] So if you multiply the both sides of this by the inverse of the left hand side's matrix...you will have your answer...why? because when you multiply a matrix by its inverse...you get the identity matrix...or just 1...leaving you with 'x = something' so what is the inverse of that matrix? \[\large \left[\begin{matrix}9 & 4 \\ 2 & 1\end{matrix}\right]\] The first step would be to find the determinant. For this 2x2 matrix it would be \( (9 \times 1) - (4 \times 2) = 9 - 8 = 1\) Great...our determinant is 1 So now that we have that...we can multiply both sides by the inverse of the matrix... We take \[\large \left[\begin{matrix}9 & 4 \\ 2 & 1\end{matrix}\right]\] switch the positions of 9 and 1 so \[\large \left[\begin{matrix}1 & 4 \\ 2 & 9\end{matrix}\right]\] and then make 2 and 4 negative \[\large \left[\begin{matrix}1 & -4 \\ -2 & 9\end{matrix}\right]\] Now we multiply this by \(\large \frac{1}{\text{determinant}}\) which we found to be 1...so just 1 So now we multiply that matrix we have now... \[\large \left[\begin{matrix}1 & -4 \\ -2 & 9\end{matrix}\right]\] to both sides of the equation \[\large \left[\begin{matrix}1 & -4 \\ -2 & 9\end{matrix}\right] \times \large \left[\begin{matrix}9 & 4 \\ 2 & 1\end{matrix}\right]x = x\] That's the left side of the equation...and for the right side \[\large \left[\begin{matrix}1 & -4 \\ -2 & 9\end{matrix}\right] \times \left[\begin{matrix}-9 & -6 \\ -1 & -8\end{matrix}\right] = \left[\begin{matrix}((1 \times -9) + (-4 \times -1)) & ((1 \times -6) + (-4 \times -8)) \\ ((-2 \times -9) + (9 \times -1)) & ((-2 \times -6) + (9 \times -8))\end{matrix}\right]\] \[\large \left[\begin{matrix}(-9 + 4)) & (-6 + 32) \\ (18 + -9) & (12 + -72)\end{matrix}\right]\] \[\large \left[\begin{matrix}-5 & 26 \\ 9 & -60\end{matrix}\right]\]

OpenStudy (anonymous):

omg thank you so much that helped a lot :)

OpenStudy (johnweldon1993):

Lol anytime :)

OpenStudy (anonymous):

could you by chance help me with one more cause this one is confusing me :P @johnweldon1993

OpenStudy (johnweldon1993):

Lol of course...same process as up there...literally gonna copy and paste this but change the numbers lol

OpenStudy (anonymous):

solve the system { 2x +y - 4z = -30 { 3x + 2y +2z =-8 {5x + 5y + z = -26 A. (-30,-8,-26) B.(4,-2,4) C. (-4,-2,4) D.(-4,2,-4)

OpenStudy (johnweldon1993):

Ahh fun! lol.... Well...want to go through the whole process? or do you know how to do it just getting thrown off?

OpenStudy (anonymous):

Yeah i would like to go through the process again because i completely forgot :P

OpenStudy (johnweldon1993):

Lol okay hang on 1 sec okay? :P

OpenStudy (johnweldon1993):

Alrighty...ready? lol

OpenStudy (kainui):

"completely forgot" just completely scroll to the top.